OpenStudy (anonymous):

Which value or values of x solve the system of equations below? y=x^2-x-6 y=x^2+6x+8 A. -2 B. -2 AND 3 C. -4 AND -2 D. -4, -2, AND 3

OpenStudy (anonymous):

equate them to each other'

OpenStudy (anonymous):

x^2 -x -6 = x^2 +6x +8 7x = -14 x=-2 when x=-2 , y= 4 +2 -6 ( by subing into the first equation ) x=-2 , y=0

OpenStudy (anonymous):

so A is the correct one

OpenStudy (anonymous):

Ooooh okay! Thank you! You really are a lifesaver.